# Activation thresholds and expressiveness of polynomial neural networks

### v1

@@ 1 Polynomial neural networks and neurovarieties
$\dim\mathcal{V}_{\mathbf{d},r}=\text{edim}\mathcal{V}_{\mathbf{d},r}\quad\text{for all}\quad r>\tilde{r}\,?$ (3)
@@
The activation threshold is the smallest $\tilde{r}$ such that (3) holds, if it exists. We denote the activation threshold of $\mathbf{d}$ by $\operatorname{ActThr}(\mathbf{d})$.
2 Expressivity for high activation degree
In this section we resolve the high activation degree conjecture of Kileel, Trager, and Bruna. Precisely, we prove [20, Conjecture 16] thereby completing the proof of [20, Theorem 14]. The key piece of our argument incorporates the number theoretic results by Newman and Slater in [26].
2.1 Powers of non-proportional polynomials
We generalize the result [26, Section 5, Equation (2)] of Newman-Slater to multivariate polynomials. The proof of this is via a simple reduction from multivariate polynomials to univariate polynomials.
Lemma 10 (Newman-Slater Corollary).
@@ 2.1 Powers of non-proportional polynomials
As the $p_{i}$ are pairwise non-proportional and the line parameterized by $\ell(t)$ is general, the univariate polynomials $p_{i}(\ell(t))$ are also pairwise non-proportional as they share no common zeros. By the result of Newman and Slater [26, Section 5, Equation (2)], we conclude that $r\leq 8(k-1)^{2}-1$. ∎
We generalize Lemma 10 further—we demonstrate that given pairwise non-proportional multivariate polynomials $p_{1},\dotsc,p_{k}$, the polynomials $p_{1}^{r},\dotsc,p_{k}^{r}$ are linearly independent for all sufficiently large $r$.
Lemma 11.
Let $\mathbb{K}$ be a subfield of $\mathbb{C}$. Given integers $d,k$ there exists an integer $\tilde{r}=\tilde{r}(k)$ with the following property. If $r>\tilde{r}$ and $p_{1},\dots,p_{k}\in\mathbb{K}[x_{1},\dots,x_{d}]$ are pairwise non-proportional, then $p_{1}^{r},\dots,p_{k}^{r}$ are linearly independent (over $\mathbb{K}$). Moreover, $\tilde{r}=8(k-1)^{2}-1$ suffices.
@@
We prove the contrapositive statement when $\mathbb{K}=\mathbb{C}$. Given integers $d,k$, let $\tilde{r}=8(k-1)^{2}-1$ and $r\geq\tilde{r}$ be an integer. Fix pairwise non-proportional polynomials $p_{1},\dotsc,p_{k}\in\mathbb{C}[x_{1},\dotsc,x_{d}]$ with the property that the set $\{p_{1}^{r},\dotsc,p_{k}^{r}\}$ is linearly dependent over $\mathbb{C}$. Then there exists $\alpha_{1},\dots,\alpha_{k}\in\mathbb{C}$ such that
@@
By Lemma 10, it follows that for some $i\neq j$, $\beta_{i}p_{i}$ and $\beta_{j}p_{j}$ are proportional. Thus, some $p_{i}$ and $p_{j}$ are proportional.
@@ 2.2 Deep networks and high activation degree
As noted in [20], Lemma 11 implies that for a fixed architecture $\mathbf{d}$, there exists $\tilde{r}$ such that for any activation degree $r>\tilde{r}$, the neurovariety $\mathcal{V}_{\mathbf{d},r}$ has the expected dimension [20, Theorem 14]. That is, Lemma 11 implies the existence of the activation threshold for any architecture $\mathbf{d}$. We recount the argument given in [20] for completeness.
@@
For fixed $\mathbf{d}=(d_{0},\dotsc,d_{L})$ satisfying $d_{i}>1$ ($i=1,\dotsc,L$), the activation threshold $\operatorname{ActThr}(\mathbf{d})$ exists. In other words, there exists $\tilde{r}$ such that whenever $r>\tilde{r}$, the neurovariety $\mathcal{V}_{\mathbf{d},r}$ has the expected dimension,
$\dim\mathcal{V}_{\mathbf{d},r}=\text{edim}\mathcal{V}_{\mathbf{d},r}=d_{L}+\sum_{i=0}^{L-1}(d_{i}-1)d_{i+1}.$
Moreover, $\operatorname{ActThr}(\mathbf{d})\leq 8\cdot(2\cdot\max\{d_{1},\dots,d_{L-1}\}-1)^{2}-1$.
@@
Fix $r>\tilde{r}=8\cdot(2\cdot\max\{d_{1},\dotsc,d_{L-1}\}-1)^{2}-1$. We show by induction on the number of layers $L$ that for activation degree $r$, a general fiber of $\Psi_{\mathbf{d},r}$ is equal to the set of multi-homogeneities of any point of the fiber given by Lemma 5. The dimension of $\mathcal{V}_{\mathbf{d},r}$ is then equal to the expected dimension.
@@
In the case that $L=1$, the neurovariety $\mathcal{V}_{\mathbf{d},r}$ has the expected dimension as the map $\Psi_{\mathbf{d},r}$ is an isomorphism.
@@
Now assume $L>1$ and fix general weights $W_{1},\dots,W_{L}$ and consider weights $\tilde{W}_{1},\dots,\tilde{W}_{L}$ such that
$(W_{L}\sigma_{L-1}W_{L-1}\dots\sigma_{1}W_{1})(x)=(\tilde{W}_{L}\sigma_{L-1}\tilde{W}_{L-1}\dots\sigma_{1}\tilde{W}_{1})(x).$ (4)
@@
Denote the output of $(W_{L-1}\sigma_{L-2}W_{L-2}\dotsb W_{2}\sigma_{1}W_{1})(x)$ by $[p_{\mathbf{w}_{1}},\dots,p_{\mathbf{w}_{d_{L-1}}}]$ and the output of $(\tilde{W}_{L-1}\sigma_{L-2}\tilde{W}_{L-2}\dotsb\tilde{W}_{2}\sigma_{1}\tilde{W}_{1})(x)$ by $[\tilde{p}_{\mathbf{w}_{1}},\dots,\tilde{p}_{\mathbf{w}_{d_{L-1}}}]$. Since the weights $W_{i}$ are general and $d_{i}>1$, the homogeneous polynomials $p_{\mathbf{w}_{i}}$ are pairwise non-proportional (there is a non-empty Zariski open set of the space of weight matrices where the outputs $p_{\mathbf{w}_{1}},\dotsc,p_{\mathbf{w}_{d_{L-1}}}$ are non-proportional). By examining the first coordinate of the outputs of (4) we obtain the decomposition
$w_{L_{11}}p_{\mathbf{w}_{1}}^{r}+\dots+w_{L_{1d_{L-1}}}p_{\mathbf{w}_{d_{L-1}}}^{r}=\tilde{w}_{L_{11}}\tilde{p}_{\mathbf{w}_{1}}^{r}+\dots+\tilde{w}_{L_{1d_{L-1}}}\tilde{p}_{\mathbf{w}_{d_{L-1}}}^{r}$ (5)
@@
where $w_{L_{ij}}$ and $\tilde{w}_{L_{ij}}$ denote the $(i,j)$ entry of the weight matrices $W_{L}$ and $\tilde{W}_{L}$ respectively. Because we have taken $r>\tilde{r}$, Lemma 11 guarantees that (5) has two proportional summands. As $p_{\mathbf{w}_{i}}$ are not pairwise proportional, no two summands on the left side may be proportional. If two summands on the right side are proportional, we reduce the number of terms on the right side by combining these terms. Otherwise, by permuting terms as necessary, we may assume these proportional terms to occur as the first term on both sides. We may then scale so that $p_{\mathbf{w}_{1}}=\tilde{p}_{\mathbf{w}_{1}}$ and subtract $\tilde{w}_{L_{11}}\tilde{p}_{\mathbf{w}_{1}}^{r}$ to obtain
$(w_{L_{11}}-\tilde{w}_{L_{11}})p^{r}_{\mathbf{w}_{1}}+\dots+w_{L_{1d_{L-1}}}p^{r}_{\mathbf{w}_{d_{L-1}}}=\tilde{w}_{L_{12}}\tilde{p}_{\mathbf{w}_{2}}^{r}+\dots+\tilde{w}_{L_{1d_{L-1}}}\tilde{p}^{r}_{\mathbf{w}d_{L-1}}.$ (6)
@@
We repeatedly apply Lemma 11 to iteratively reduce the number of terms on the right side, until the right side of (6) is zero. Then, the pairwise linear independence of the $p_{\mathbf{w}_{i}}$ and Lemma 11 imply that each coefficient is also zero. Hence, up to scaling and permutation, $[p_{\mathbf{w}_{1}},\dots,p_{\mathbf{w}_{d_{L-1}}}]=[\tilde{p}_{\mathbf{w}_{1}},\dots,\tilde{p}_{\mathbf{w}_{d_{L-1}}}]$ and the entries of $W_{L}$ and $\tilde{W}_{L}$ agree in the first row. This process is repeated for each row of $W_{L}$ and $\tilde{W}_{L}$ to show that they are equal (up to scaling and permutation). Therefore the fiber $\Psi_{\mathbf{d},r}^{-1}(\Psi_{\mathbf{d},r}(W_{1},\dots,W_{L}))$ has dimension $\sum_{i=1}^{L-1}d_{i}.$

### v3

@@ 1.2 Main result and example
$\dim\mathcal{V}_{\mathbf{d},r}=\mathrm{edim}\mathcal{V}_{\mathbf{d},r}\quad\text{for all}\quad r>\tilde{r}\,?$ (5)
When such a non-negative integer $\tilde{r}$ exists for (5) we have the following definition.
@@
The activation degree threshold of an architecture $\mathbf{d}=(d_{0},d_{1},\dots,d_{L})$ with $L>1$, is
$\operatorname{ActThr}(\mathbf{d}):=\min\{\tilde{r}\in\mathbb{N}_{\geq 0}:\dim\mathcal{V}_{\mathbf{d},r}=\mathrm{edim}\mathcal{V}_{\mathbf{d},r}\text{ for all }r>\tilde{r}\}.$
@@
In other words, $\operatorname{ActThr}(\mathbf{d})$ is the smallest non-negative integer $\tilde{r}$ such that (5) holds.
With this terminology in hand, we state our main result, which shows the activation degree threshold is well defined when $\mathbf{d}$ has no entries equal to one.
Theorem 18.
For fixed $\mathbf{d}=(d_{0},d_{1},\dotsc,d_{L})$ satisfying $d_{i}>1$ ($i=0,\dotsc,L-1$), the activation degree threshold $\operatorname{ActThr}(\mathbf{d})$ exists. In other words, there exists an integer $\tilde{r}(\mathbf{d})$ such that whenever $r>\tilde{r}(\mathbf{d})$, the neurovariety $\mathcal{V}_{\mathbf{d},r}$ of the polynomial neural network has the expected dimension,
$\dim\mathcal{V}_{\mathbf{d},r}=\mathrm{edim}\mathcal{V}_{\mathbf{d},r}=d_{L}+\sum_{i=0}^{L-1}(d_{i}-1)d_{i+1}.$
Moreover, for $L>1$, we have
$\operatorname{ActThr}(\mathbf{d})\leq 6{m}^{2}-6m+1,\quad m:=-1+2\max\{d_{1},\dots,d_{L-1}\}.$ (6)
Example 11.
This example aims to illustrate the activation degree threshold by mapping the function space of neural networks with architecture $(3,2,1)$ and activation function $x^{r}$ for $r\in\mathbb{R}_{>0}$. To ensure these functions are well defined, we restrict $\mathbb{R}^{d_{0}}$ to the nonnegative orthant and assume that each weight matrix has nonnegative entries. Denote the corresponding parameter space and functional space as $\Theta^{+}$ and $\operatorname{Fun}^{+}_{r}\left(\mathbb{R}^{3},\mathbb{R}^{1}\right)$ respectively.
Note that each point $\mathbf{x}$ in $\mathbb{R}^{3}$ induces the map $E_{\mathbf{x}}:\operatorname{Fun}_{r}^{+}\left(\mathbb{R}^{3},\mathbb{R}^{1}\right)\to\mathbb{R}^{1}$ by evaluation $\psi\mapsto\psi(\mathbf{x})$. The neuromanifold $\operatorname{Fun}^{+}_{r}\left(\mathbb{R}^{3},\mathbb{R}^{1}\right)$ is known to have dimension at most $6$ for all $r$. Therefore, consider the six points $\mathbf{x}_{1},\dots,\mathbf{x}_{6}\in\mathbb{R}^{d_{0}}$ so that $(E_{\mathbf{x}_{1}},\dots,E_{\mathbf{x}_{6}})$ induces a linear map from $\operatorname{Fun}_{r}^{+}(\mathbb{R}_{\geq 0}^{3},\mathbb{R}_{\geq 0})$ to $\mathbb{R}_{\geq 0}^{6}$. Precomposing with the parameter map (1) we get a family of functions depending on $r$
$\mathbb{R}^{3\times 2}\times\mathbb{R}^{2\times 1}{\supset\Theta^{+}}\to{E\left(\operatorname{Fun}_{r}^{+}(\mathbb{R}_{\geq 0}^{3},\mathbb{R}_{\geq 0})\right)}\subset\mathbb{R}^{6},\quad\theta=(W_{1},W_{2})\mapsto E\left(F_{\theta}\right).$ (7)
The benefit of this setup is that it allows for a straightforward rank computation of the Jacobian matrix. For example, we calculate the Jacobian matrix of the map (7) with the following choices. The six points we choose to define $E$ are $(1,0,0)$, $(0,1,0)$, $(0,0,1)$, $(1,1,0)$, $(1,0,1)$, $(0,1,1)$, and we calculate the Jacobian of (7) at the point
$\left(\left[\begin{array}[]{cc}1&1\\ 1&2\\ 1&3\end{array}\right],\left[\begin{array}[]{c}1\\ 1\end{array}\right]\right)\in\mathbb{R}^{3\times 2}\times\mathbb{R}^{2\times 1}.$
The resulting Jacobian matrix is
[$r&0&0&r&0&0&1&1\\ 0&r&0&0&r2^{r-1}&0&1&2^r\\ 0&0&r&0&0&r3^{r-1}&1&3^r\\ r2^{r-1}&r2^{r-1}&0&r3^{r-1}&r3^{r-1}&0&2^r&3^r\\ r2^{r-1}&0&r2^{r-1}&r4^{r-1}&0&r4^{r-1}&2^r&4^r\\ 0&r2^{r-1}&r2^{r-1}&0&r5^{r-1}&r5^{r-1}&2^r&5^r\\ \end{array}\right].$ (8)
Denote the $6\times 6$ submatrix on the left by $J(r)$. The positive integers where the $J(r)$ drops rank are $r=1,2$. Therefore the activation degree threshold of this architecture is less than or equal to three.
We visualize the functions $\det(J(r)$ and $\log|\det(J(r))|$ in Figure 1. Explicitly, the value of $\det(J(r))$ is the following polynomial exponential:
$r^{6}\left(-96^{r-1}-32^{r-1}+3\cdot 48^{r-1}+80^{r-1}-36^{r-1}-3\cdot 60^{r-1}+90^{r-1}+30^{r-1}\right).$ (9)
Figure 1: The blue curve is the graph of $r\mapsto\frac{1}{r^{6}}\det(J(r))$ where $J(r)$ is the $6\times 6$ submatrix of (8). The red curve is $(r,\log|y|)$ where $(r,y)$ is on the blue curve. The orange dashed lines are $r=1$ and $r=2$.
Remark 12.
For some intuition, we relate the activation degree threshold to $\operatorname{defect}(\mathcal{V}_{\mathbf{d},r})$. For fixed $\mathbf{d}=(d_{0},d_{1},\dotsc,d_{L})$ satisfying $d_{i}>1$ ($i=0,\dotsc,L-1$),
$\operatorname{defect}(\mathcal{V}_{\mathbf{d},r})=0,\text{ for all }r>\operatorname{ActThr}(\mathbf{d}),$
and $\operatorname{ActThr}(\mathbf{d})=0$ or $\operatorname{defect}(\mathcal{V}_{\mathbf{d},\operatorname{ActThr}(\mathbf{d})})$ is positive.
2 Expressiveness for high activation degree
In this section we develop results on the linear independence of powers of polynomials (Section 2.1 and Section 2.2) to prove our main result in Section 2.3. In particular, our Theorem 18 completes the results in Corollary 17.
2.1 A prior number theoretic result by Newman-Slater
In this section, we recover a number theoretic result by Newman-Slater in [34]. The proof here slightly improves their bound.
Theorem 13.
Let $n>1$ and
$R_{1}^{n}(x)+R_{2}^{n}(x)+\cdots+R_{k}^{n}(x)=1$
where $R_{1},R_{2},\dots,R_{k}$ are non-constant rational functions in the variable $x$. Then $n\leq 6{k}^{2}-6k+1$.
Proof.
A proof of this statement follows from modifications of the proof of [34, Section 4, Theorem]. We include it here for the benefit of the reader.
Consider the $k\times k$ Wronskian matrix $W:=W(R_{1}^{n},\dots,R_{k}^{n})$ where the $(i,j)$-th entry is the $(i-1)$-st derivative of $R_{j}^{n}$, which we denote by $\partial^{i-1}R_{j}^{n}$. With this setup, we have the following vector equation, with the first standard basis vector on the right hand side and $\mathbf{1}$ denoting the vector of ones:
$W\cdot\mathbf{1}=\mathbf{e}_{1}.$
By letting $D$ be the $k\times k$ diagonal matrix with $(j,j)$ entry ${R_{j}^{-n+k-1}}$ we have
$\left(W\cdot D\right)\left(D^{-1}\cdot\mathbf{1}\right)=\mathbf{e}_{1}$
where
$W\cdot D=\begin{bmatrix}R_{1}^{k-1}&\dots&R_{k}^{k-1}\\ \vdots&\dots&\vdots\\ \dfrac{\partial^{k-1}R_{1}^{n}}{R_{1}^{n-k+1}}&\dots&\dfrac{\partial^{k-1}R_{k}^{n}}{R_{k}^{n-k+1}}\\ \end{bmatrix}.$
The benefit of this formulation is that cancellation of common factors in numerators and denominators occurs with the multiplication $W\cdot D$.
If the rational functions $R_{1}^{n},\dotsc,R_{k}^{n}$ are linearly independent, then $\det(W)\neq 0$ and hence $\det(W\cdot D)\neq 0$. Using Cramer’s rule, we solve for the first entry of $\left(D^{-1}\cdot\mathbf{1}\right)$:
$R_{1}^{n-k+1}=\det(B)/\det(W\cdot D),$ (10)
where $B$ is obtained by replacing the first column of $W\cdot D$ with $\mathbf{e}_{1}$.
Assume each $R_{j}$ is written in lowest terms. After reordering the $R_{j}$’s as necessary, suppose $R_{1}$ has the largest degree numerator or denominator and denote this degree by $A$. Hence, the sum of the degrees of the numerator and denominator on the left hand side of (10) is between $A\cdot|n-k+1|$ and $2A\cdot|n-k+1|$.
Now note that if $f$, $g$ are polynomials, then
$(f/g^{r})^{\prime}=(f^{\prime}g^{r}-frg^{\prime}g^{r-1})/g^{2r}=(f^{\prime}g-frg^{\prime})/g^{r+1}.$
In other words, the sum of the degrees of the numerator and denominator, when taking the derivative once, increases by no more than $12\deg(g)-$ 1. As a consequence, the $(i,j)$-th entry of the matrix $W\cdot D$ when written in lowest terms is a rational function for which the sum of the degrees of the numerator and denominator is no more than $2A(k-1)+(2A-1)(i-1)$.
By the Leibniz formula for determinant, the sum of the degrees of the numerator and denominator for $\det(B)$ is no more than
$\sum_{i=2}^{k}\big{(}2A\cdot(k-1)+(2A-1)(i-1)\big{)}=$ $2\displaystyle 2A\cdot(k-1)^{2}\,+\,(2A-1)k(k-1)/$ 2
$=$ $A\cdot(k-1)(3k-2)\,-\,k(k-1)/2$
while the sum of the degrees of the numerator and denominator for $\det(W\cdot D)$ is no more than
$\sum_{i=1}^{k}\big{(}2A\cdot(k-1)+(2A-1)(i-1)\big{)}=$ $2\displaystyle 2A\cdot k(k-1)+(2A-1)k(k-1)/$ 2
$=$ $A\cdot(k-1)(3k)-k(k-1)/2.$
So we have
${\color[rgb]{0,0,0}A\cdot(n-k{+1})\leq A\cdot|n-k+1|}\leq\Big{(}A\cdot(k-1)(3k-2)-k(k-1)/2\Big{)}+\Big{(}A\cdot(k-1)(3k)-k(k-1)/2\Big{)}.$
This implies the inequality
$n\leq\left(6-\frac{1}{A}\right)k^{2}+\left(-7+\frac{1}{A}\right)k+1.$
As $A\geq 1$, we arrive at
$n\leq 6{k}^{2}-6k+1.$ (11)
We have shown the inequality (11) holds when $R_{1}^{n},\dots,R_{k}^{n}$ are linearly independent. If these functions are linearly dependent, then we may find a subset $S$ of $\{1,\dots,k\}$ with cardinality $\ell$ less than $k$ such that $\sum_{j\in S}R_{j}^{n}=1$ and $\{R_{j}^{n}:j\in S\}$ is linearly independent. Applying the Wronskian argument in this setting proves
$n\leq 6{\ell}^{2}-6\ell+1.$
Since $1<\ell<k$ implies $16{\ell}^{2}-6\ell+1<6{k}^{2}-6k+$ 1, the theorem holds.
Corollary 14.
Let $P_{1},\dots,P_{k}$ and $R$ be polynomials in the variable $x$ such that $R$ is not the zero polynomial and no $P_{i}$ is proportional to $R$. If $P_{1}^{n}+P_{2}^{n}+\cdots+P_{k}^{n}=R^{n}$, then $n\leq 6{k}^{2}-6k+1$.
Proof.
Since $R$ is not the zero polynomial, the equation $P_{1}^{n}+P_{2}^{n}+\cdots+P_{k}^{n}=R^{n}$ holds if and only if
$\left(\frac{P_{1}}{R}\right)^{n}+\left(\frac{P_{2}}{R}\right)^{n}+\cdots+\left(\frac{P_{k}}{R}\right)^{n}=1.$
The result follows from Theorem 13 since each $P_{i}/R$ is a non-constant rational function. ∎
2.2 Powers of non-proportional multivariate polynomials
The statements in the previous subsection are for univariate polynomial functions. In this subsection, we extend those results to the multivariate polynomial setting to apply them to polynomial neural networks in Section 2.3.
To that end, first we generalize Corollary 14 to multivariate polynomials.
Lemma 15.
@@ 2.2 Powers of non-proportional multivariate polynomials
As the $p_{i}$ are pairwise non-proportional and the line parameterized by $\ell(t)$ is general, the univariate polynomials $p_{i}(\ell(t))$ are also pairwise non-proportional as they share no common zeros. By Corollary 14, we conclude that
$r\leq 6{(k-1)}^{2}-6(k-1)+1.$
Next we demonstrate that given pairwise non-proportional multivariate polynomials $p_{1},\dotsc,p_{k}$, the polynomials $p_{1}^{r},\dotsc,p_{k}^{r}$ are linearly independent for all sufficiently large $r$.
Proposition 16.
Let $\mathbb{K}$ be a subfield of $\mathbb{C}$. Given integers $d,k$, there exists an integer $\tilde{r}=\tilde{r}(k)$ with the following property. If $r>\tilde{r}(k)$ and $p_{1},\dots,p_{k}\in\mathbb{K}[x_{1},\dots,x_{d}]$ are pairwise non-proportional, then $p_{1}^{r},\dots,p_{k}^{r}$ are linearly independent (over $\mathbb{K}$). Moreover, $\tilde{r}(k)=6{(k-1)}^{2}-6(k-1)+1$ has the desired property.
@@
We prove the contrapositive statement when $\mathbb{K}=\mathbb{C}$. Given integers $d,k$, let $\tilde{r}(k)=6{(k-1)}^{2}-6(k-1)+1$ and $r\geq\tilde{r}(k)$ be an integer. Fix pairwise non-proportional polynomials $p_{1},\dotsc,p_{k}\in\mathbb{C}[x_{1},\dotsc,x_{d}]$ with the property that the set $\{p_{1}^{r},\dotsc,p_{k}^{r}\}$ is linearly dependent over $\mathbb{C}$. Then there exist $\alpha_{1},\dots,\alpha_{k}\in\mathbb{C}$ such that
@@
By Lemma 15, it follows that for some $i\neq j$, $\beta_{i}p_{i}$ and $\beta_{j}p_{j}$ are proportional. Thus, some $p_{i}$ and $p_{j}$ are proportional. ∎
Corollary 17 (Conjecture 16, [25]).
There exists $\tilde{r}:\mathbb{N}^{2}_{>0}\to\mathbb{Z}$, $(d,k)\mapsto\tilde{r}(d,k)$ with the following property. Whenever $p_{1},...,p_{k}\in\mathbb{R}[x_{1},...,x_{d}]$ are $k$ homogeneous polynomials of the same degree in $d$ variables, no two of which are linearly dependent, $p^{r}_{1},\dots,p^{r}_{k}$ are linearly independent if $r>\tilde{r}({d,k})$.
Proof.
Note that $p_{i},p_{j}$ being linearly independent implies $p_{i},p_{j}$ are non-proportional. By Proposition 16 it follows that $\tilde{r}(d,k)=6{(k-1)}^{2}-6(k-1)+1$ suffices. ∎
@@ 2.3 Deep networks and high activation degree
Next, we show when activation degree thresholds exist for polynomial neural networks. The proof of the following theorem follows the arguments for [25, Theorem 14], except where we leverage our Proposition 16.
@@
For fixed $\mathbf{d}=(d_{0},d_{1},\dotsc,d_{L})$ satisfying $d_{i}>1$ ($i=0,\dotsc,L-1$), the activation degree threshold $\operatorname{ActThr}(\mathbf{d})$ exists. In other words, there exists an integer $\tilde{r}(\mathbf{d})$ such that whenever $r>\tilde{r}(\mathbf{d})$, the neurovariety $\mathcal{V}_{\mathbf{d},r}$ of the polynomial neural network has the expected dimension,
$\dim\mathcal{V}_{\mathbf{d},r}=\mathrm{edim}\mathcal{V}_{\mathbf{d},r}=d_{L}+\sum_{i=0}^{L-1}(d_{i}-1)d_{i+1}.$
Moreover, for $L>1$, we have
$\operatorname{ActThr}(\mathbf{d})\leq 6{m}^{2}-6m+1,\quad m:=-1+2\max\{d_{1},\dots,d_{L-1}\}.$ (12)
@@
Fix $r>\tilde{r}(\mathbf{d})=6{m}^{2}-6m+1$. The expected dimension of a neurovariety $\mathcal{V}_{\mathbf{d},r}$ is the minimum of two numbers in (4). We may assume
$\operatorname{edim}(\mathcal{V}_{\mathbf{d},r})=d_{L}+\sum_{i=0}^{L-1}(d_{i}d_{i+1}-d_{i+1})\leq d_{L}\binom{d_{0}+r^{L-1}-1}{r^{L-1}}.$
When $L=1$ and $r\geq 1$, this can be verified by direct computation. If $L>1$ then $d_{L}\binom{d_{0}+r^{L-1}-1}{r^{L-1}}$ is no less than $12\binom{2+r^{L-1}-1}{r^{L-1}}=2r^{L-1}+$ 1, while $d_{L}+\sum_{i=0}^{L-1}(d_{i}d_{i+1}-d_{i+1})$ is no more than $(L-1)B^{2}$ with $B:=\max\{d_{1},\dots,d_{L-1}\}$.
Thus we have
$r^{L-1}<2r^{L-1}+1<(L-1)B^{2}$
which implies that
$r<(L-1)^{1/(L-1)}B^{2/(L-1)}<2B^{2}<\tilde{r}(\mathbf{d}).$
We show by induction on the number of layers that for activation degree $r$, a general fiber of $\Psi_{\mathbf{d},r}$ is equal to the set of multi-homogeneities of any point of the fiber given by Lemma 4. The dimension of $\mathcal{V}_{\mathbf{d},r}$ is then equal to the expected dimension.
In the case that $L=1$, the neurovariety $\mathcal{V}_{\mathbf{d},r}$ has the expected dimension as the map $\Psi_{\mathbf{d},r}$ is an isomorphism. Therefore, the activation degree threshold for the architecture $(d_{0},d_{1})$ is zero and satisfies (12).
@@
Now assume $L>1$. Fix general weights $W_{1},W_{2},\dots,W_{L}$ and consider given weights $\tilde{W}_{1},\tilde{W}_{2}\dots,\tilde{W}_{L}$ such that
$\begin{split}&(W_{L}\sigma_{L-1}W_{L-1}\dots\sigma_{1}W_{1})(x)\\ &=(\tilde{W}_{L}\sigma_{L-1}\tilde{W}_{L-1}\dots\sigma_{1}\tilde{W}_{1})(x).\end{split}$ (13)
Denote the output of
$(W_{L-1}\sigma_{L-2}W_{L-2}\dotsb W_{2}\sigma_{1}W_{1})(x)$
by $[p_{\mathbf{w}_{1}},\dots,p_{\mathbf{w}_{d_{L-1}}}]$ and the output of
$(\tilde{W}_{L-1}\sigma_{L-2}\tilde{W}_{L-2}\dotsb\tilde{W}_{2}\sigma_{1}\tilde{W}_{1})(x)$
@@
by $[\tilde{p}_{\mathbf{w}_{1}},\dots,\tilde{p}_{\mathbf{w}_{d_{L-1}}}]$. Since the weights $W_{i}$ are general and $d_{i}>1$, the homogeneous polynomials $p_{\mathbf{w}_{i}}$ are pairwise non-proportional (there is a non-empty Zariski open set of the space of weight matrices where the outputs $p_{\mathbf{w}_{1}},\dotsc,p_{\mathbf{w}_{d_{L-1}}}$ are non-proportional). By examining the first coordinate of the outputs of (13) we obtain the decomposition
$\begin{split}&w_{L_{11}}p_{\mathbf{w}_{1}}^{r}+{{w}_{L_{12}}{p}_{\mathbf{w}_{2}}^{r}}+\dots+w_{L_{1d_{L-1}}}p_{\mathbf{w}_{d_{L-1}}}^{r}\\ &=\tilde{w}_{L_{11}}\tilde{p}_{\mathbf{w}_{1}}^{r}+{\tilde{w}_{L_{12}}\tilde{p}_{\mathbf{w}_{2}}^{r}}+\dots+\tilde{w}_{L_{1d_{L-1}}}\tilde{p}_{\mathbf{w}_{d_{L-1}}}^{r}\end{split}$ (14)
@@
where $w_{L_{ij}}$ and $\tilde{w}_{L_{ij}}$ denote the $(i,j)$ entry of the weight matrices $W_{L}$ and $\tilde{W}_{L}$ respectively. Because we have taken $r>\tilde{r}$, Proposition 16 guarantees that (14) has two proportional summands. As $p_{\mathbf{w}_{i}}$ are not pairwise proportional, no two summands on the right side may be proportional. The equality implies that neither may two summands on the left side be proportional.
By permuting terms as necessary, we may assume these proportional terms to occur as the first term on both sides. We may then scale so that $p_{\mathbf{w}_{1}}=\tilde{p}_{\mathbf{w}_{1}}$ and subtract $\tilde{w}_{L_{11}}\tilde{p}_{\mathbf{w}_{1}}^{r}$ to obtain
$\begin{split}(w_{L_{11}}-\tilde{w}_{L_{11}})p^{r}_{\mathbf{w}_{1}}+{{w}_{L_{12}}{p}_{\mathbf{w}_{2}}^{r}}+\dots+w_{L_{1d_{L-1}}}p^{r}_{\mathbf{w}_{d_{L-1}}}\\ =\tilde{w}_{L_{12}}\tilde{p}_{\mathbf{w}_{2}}^{r}+\dots+\tilde{w}_{L_{1d_{L-1}}}\tilde{p}^{r}_{\mathbf{w}d_{L-1}}.\end{split}$ (15)
@@
We repeatedly apply Proposition 16 to iteratively reduce the number of terms on the right side, until the right side of (15) is zero. Then the pairwise linear independence of the $p_{\mathbf{w}_{i}}$ and Proposition 16 imply that each coefficient is also zero. Hence, up to scaling and permutation, $[p_{\mathbf{w}_{1}},\dots,p_{\mathbf{w}_{d_{L-1}}}]=[\tilde{p}_{\mathbf{w}_{1}},\dots,\tilde{p}_{\mathbf{w}_{d_{L-1}}}]$ and the entries of $W_{L}$ and $\tilde{W}_{L}$ agree in the first row. This process is repeated for each row of $W_{L}$ and $\tilde{W}_{L}$ to show that they are equal (up to scaling and permutation). Therefore the fiber $\Psi_{\mathbf{d},r}^{-1}(\Psi_{\mathbf{d},r}(W_{1},W_{2},\dots,W_{L}))$ has dimension $\sum_{i=1}^{L-1}d_{i}.$
Corollary 19 (Conjecture 5.2, [28]).
For any fixed widths $\mathbf{d}=(d_{0},\dots,d_{L})$ with $d_{i}>1$ for $i=1,\dots,L-1$, there exists an integer $\tilde{r}(\mathbf{d})$ such that whenever $r>\tilde{r}(\mathbf{d})$, the neurovariety $\mathcal{V}_{\mathbf{d},r}$ of a polynomial neural network attains the expected dimension.
Table 1 of [28] provides several examples of neurovarieties that have the expected dimension. In this next section we show that equi-width architectures have activation degree threshold one, that is, the corresponding neurovarieties always have the expected dimension when the activation degree is two or more.
Remark 20.
In the proof of Theorem 18, we leveraged a surprising result about pairwise non-proportionality implying linear independence. Linear independence of powers of polynomials is part of a larger research program of understanding Fermat hypersurfaces and tickets [38] of polynomial systems.